I think I am doing something wrong maybe someone can tell me what Did wrong.
The question is Brian starts walking from point A at 2 mph. One-half hour later, Elliott starts walking from point A 3 1/2 mph follows the same route.
How long will it take Elliott to catch up with Brian?
This is what I did
RATE x TIME = DISTANCE
BRIAN_ 2 x 2x
ELLIOTT 3.5 1.5 3.5 (1.5)
2x=3.5(1.5)
2x=5.25
x= 2.62
is this right?????

Answer by stanbon(48551)

Brian starts walking from point A at 2 mph. One-half hour later, Elliott starts walking from point A 3 1/2 mph follows the same route.
How long will it take Elliott to catch up with Brian?
Brian DATA:
rate=2mph; time= x hours; distance= 2x miles
Elliott DATA:
rate=(7/2)mph; time= x-(1/2) hrs.; distance = (7/2)(x-1/2) miles
EQUATION:
distance = distance
2x=(7/2)x- (7/4)
(8/4)x = 14/4 x = 7/4
-6/4 x = - 7/4
x=7/6 hours = 1 1/6 hours = 1 hr 10min
Cheers,
Stan H.